# Generalized Random Forests using Fixed-Point Trees

### v3

@@ 4 Theoretical Analysis
Under the conditions of Theorem 4.2, suppose moreover that Regularity Condition 1 holds, and that a forest is grown on subsamples of size $s$ scaling as $s=n^{\beta}$, satisfying Regularity Condition 2. Then, there exists a sequence $\sigma_{n}(x)$ such that $(\hat{\theta}_{n}(x)-\theta^{*}(x))/\sigma_{n}(x)\leadsto\mathcal{N}(0,1)$ and $\sigma^{2}_{n}(x)=\textnormal{polylog}(n/s)^{-1}s/n$, where $\textnormal{polylog}(n/s)$ is a function that is bounded away from 0 and increases at most polynomially with the log of the inverse sampling ratio $\log(n/s)$.
The proof of Theorem 4.3 an immediate consequence of Theorem 5 of Athey et al. (2019). Theorems 4.2 and 4.3 demonstrate that the improved computational efficiency of the GRF-FPT algorithm does not compromise its key asymptotic properties. Specifically, standard errors and confidence intervals derived for GRF estimates remain valid for GRF-FPT estimates $\hat{\theta}(x)$.
@@ 5 Applications
where $\nu^{*}(\cdot)$ is a nuisance intercept function, $\theta^{*}(x)=(\theta^{*}_{1}(x),\ldots,\theta^{*}_{K}(x))^{\top}$ are the target effect functions local to $X_{i}=x$, under the assumptions $\mathbb{E}[\epsilon_{i}|X_{i}=x]=0$ and $\mathbb{E}[\epsilon_{i}W_{i}|X_{i}=x]={\bf 0}$.
Varying coefficient models (VCM). Given regressors $W_{i}\in\mathbb{R}^{K}$, models of the form (25) can be characterized as varying coefficient models (Hastie & Tibshirani, 1993). As discussed in Section 3.2, we must also assume that the regressors $W_{i}$ are conditionally exogeneous given $X_{i}=x$.
@@
$\psi_{\theta,\nu}(Y_{i},W_{i})\coloneqq\begin{bmatrix}Y_{i}-W_{i}^{\top}\theta-\nu\\ (Y_{i}-W_{i}^{\top}\theta-\nu)W_{i}\end{bmatrix}.$
@@
$A_{P}=-\frac{1}{n_{P}}\sum_{\{i:X_{i}\in P\}}(W_{i}-\overline{W}_{P})(W_{i}-\overline{W}_{P})^{\top}.$
@@
$\hat{\theta}_{P}\coloneqq-A_{P}^{-1}\frac{1}{n_{P}}\sum_{\{i:X_{i}\in P\}}(W_{i}-\overline{W}_{P})(Y_{i}-\overline{Y}_{P}).$
@@
$\rho^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{i}$ $=-\xi^{\top}\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(Y_{i},W_{i}),$
@@
The relationship $\rho_{i}^{\texttt{grad}}=A_{P}^{-1}\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ reveals a significant benefit of fixed-point pseudo-outcomes. The form of $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ eliminates the computational cost associated with the multiplication of $A_{P}^{-1}$. Furthermore, the computation of $\hat{\theta}_{P}$ in (27) no longer requires solving for $A_{P}^{-1}$. Therefore, to further enhance computational efficiency, we consider a one-step gradient descent approximation $\tilde{\theta}_{P}$ of the true OLS coefficients $\hat{\theta}_{P}$, taken from the origin:
$\tilde{\theta}_{P}\coloneqq\gamma\frac{1}{n_{P}}\sum_{\{i:X_{i}\in P\}}(W_{i}-\overline{W}_{P})(Y_{i}-\overline{Y}_{P}),$ (28)
where $\gamma$ denotes the exact line search step size for the regression of $Y_{i}-\overline{Y}_{P}$ on $W_{i}-\overline{W}_{P}$ over $P$:
@@
where $W=\left[\begin{matrix}W_{1}~{}\cdots~{}W_{n_{P}}\end{matrix}\right]^{\top}$ and $Y=\left[\begin{matrix}Y_{1}~{}\cdots~{}Y_{n_{P}}\end{matrix}\right]^{\top}$ with the centering $W-\overline{W}_{P}$ and $Y-\overline{Y}_{P}$ understood as applied row-wise.
The computational cost associated with $\tilde{\theta}_{P}$ is small because many of the products appearing in (28) and (29) are pre-computed as part of $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ in (27). These accelerations are particularly compelling when the dimension of $\theta^{*}(x)$ is large and computational efficiency is critical, as in large-scale A/B testing with multiple concurrent treatment arms or observational studies with numerous treatment levels (Kohavi et al., 2013; Bakshy et al., 2014).
@@ 6 Simulations
In this section, we empirically compare the computational efficiency and estimation accuracy of the accelerated fixed-point method, GRF-FPT, and the gradient-based method, GRF-grad, under the VCM and HTE designs. Implementation details and links to the reproducible code are provided in Appendix C.4.
Settings. We follow the structural model in (25), where auxiliary variables are generated as $X_{i}\sim\mathcal{U}([0,1]^{p})$ with noise $\epsilon_{i}\sim\mathcal{N}(0,1)$. The outcome $Y_{i}$ follows (25). For VCM experiments, regressors $W_{i}\in\mathbb{R}^{K}$ are sampled from $\mathcal{N}_{K}({\bm{0}},\mathbb{I})$, while for HTE experiments, $W_{i}\in\{0,1\}^{K}$ follows a multinomial distribution, $W_{i}\,|\,X_{i}=x\sim\text{Multinomial}(1,(\pi_{1}(x),\ldots,\pi_{K}(x)))$, where $\pi_{k}(x)$ is the probability of treatment level $k\in\{1,\ldots,K\}$. We set $\nu^{*}(x)\coloneqq 0$ and vary the target effect functions $\theta^{*}_{k}(x)$ and treatment probabilities $\pi_{k}(x)$ across different settings, fully detailed in Appendix C.4.
Results. To evaluate computational efficiency, we compare training times for forests of 1, 50, and 100 trees. Figures 6-9 in Appendix D show that GRF-FPT consistently achieves a 2-3x speedup over GRF-grad, with efficiency gains increasing as the target dimension $K$ grows. This trend persists in large-scale experiments with up to 1000 trees, as shown in Figures 2 and 3, where GRF-FPT maintains a 2-2.5x speed advantage. Additional timing results for VCM and HTE experiments are presented in Appendix D.
Figure 2: Absolute fit time for VCM experiments for $K=256$-dimensional targets.
To assess estimation accuracy, we evaluate the mean squared error (MSE) of $\hat{\theta}(x)$ across 40 replications, each training a forest of 100 trees on $n=20,000$ observations and testing on a separate set of $5,000$ observations. Figure 10 in Appendix D confirms that GRF-FPT matches the accuracy of GRF-grad while significantly reducing computation time. Further comparisons for both VCM and HTE settings are provided in Appendix D.
Figure 3: Speedup factor for VCM experiments for $K=256$-dimensional targets.
@@ 7 Real Data Application
Data. In this section we apply GRF-FPT to the analysis of geographically-varying effects $\theta^{*}(x)$ on housing prices. The data, first appearing in Kelley Pace & Barry (1997), contains 20,640 observations of housing prices taken from the 1990 California census. Each observation corresponds to measurements aggregated over a small geographical census block and contains measurements of 9 variables: median housing value, longitude, latitude, median housing age, total rooms, total bedrooms, population, households, and median income. We employ a varying coefficient model of the form (25) where $Y_{i}$ is the housing value, $X_{i}$ are spatial coordinates, and $W_{i}=(W_{i,1},\ldots,W_{i,6})^{\top}$ are the remaining six regressors. Details of the model and data transformations used for the California housing analysis is found in Appendix F.
Results. Table 4 summarizes the computational benefit of GRF-FPT applied to the California housing data. Figure 5 illustrates the six geographically-varying effect estimates under GRF-FPT, with qualitatively similar results shown in Figure 20 for GRF-grad in Appendix F.
Figure 5: Geographically-varying GRF-FPT estimates $\hat{\theta}(x)$.
@@ 8 Conclusion
Our results demonstrate that the fixed-point tree algorithm offers a substantial computational advantage over the gradient-based approach, particularly in large-dimensional settings where savings scale as $\mathcal{O}(K^{2})$, without compromising in accuracy. This highlights GRF-FPT as a powerful alternative for large-scaled estimation problems, particularly for problems where estimates of the target function must be learned from the data rather than observed directly. Future work may explore extensions to larger-scale problems and alternative function estimation tasks, such as unsupervised learning and structured prediction. Our findings position GRF-FPT as a scalable and robust alternative for practitioners seeking efficient localized estimation methods.
9 Impact Statement
@@ A.1 Assumptions
We follow the key assumptions of Athey et al. (2019). In our theoretical analyses, the predictor and parameter spaces are both subsets of Euclidean space such that $x\in\mathcal{X}=[0,1]^{p}$ and $(\theta,\nu)\in\mathcal{B}\subset\mathbb{R}^{K}$, where $\mathcal{B}$ is a compact subset of $\mathbb{R}^{K}$. Under the analyses of Wager & Walther (2015), we suppose that the features of the covariates $X_{i}=(X_{i,1},\ldots,X_{i,p})^{\top}$ have density $f_{X}$ that is bounded away from 0 and $\infty$, i.e. $c\leq f_{X}(x)\leq C<\infty$, for some constants $c>0$ and $C<\infty$. We do not require that the score function $\psi$ be continuous in $(\theta,\nu)$, as is the case for quantile estimation, but we do require that the expected score function
$M_{\theta,\nu}(x)\coloneqq\mathbb{E}_{O|X}\left[\psi_{\theta,\nu}(O_{i})\middle|X_{i}=x\right],$ (30)
is smooth with respect to its parameters $(\theta,\nu)$.
@@
    The score functions $\psi_{\theta,\nu}(O_{i})$ have a continuous covariance structure in the following sense: Let $\gamma(\cdot,\cdot)$ denote the worst-case variogram,
@@
    The estimating $\psi$-function can be written as
@@
    for $\lambda$ a Lipschitz-continuous function in $(\theta,\nu)$, $g:\{O_{i}\}\to\mathbb{R}$ a univariate summary of the observables $O_{i}$, and $\zeta_{\theta}:\mathbb{R}\to\mathbb{R}$ any family of monotone and bounded functions.
@@
    For any weights $\alpha_{i}$, $\sum\alpha_{i}=1$, the minimizer $(\hat{\theta},\hat{\nu})$ of the weighted empirical estimation problem (5) is at least an approximate root of the objective function $\sum^{n}_{i=1}\alpha_{i}\psi_{\theta,\nu}(O_{i})$,
@@
    The estimating function $\psi_{\theta,\nu}(O_{i})$ is a negative sub-gradient of a convex function, and the expected estimating function $M_{\theta,\nu}(X_{i})$ is the negative gradient of a strongly convex function.

### v4

@@ 4 Theoretical Analysis
Under the conditions of Theorem 4.3, suppose moreover that Regularity Condition 1 holds, and that a forest is grown on subsamples of size $s$ scaling as $s=n^{\beta}$, where $\beta$ satisfies Regularity Condition 2. Then, there exists a sequence $\sigma_{n}(x)$ such that $(\hat{\theta}_{n}(x)-\theta^{*}(x))/\sigma_{n}(x)\leadsto\mathcal{N}(0,1)$ and $\sigma^{2}_{n}(x)=\textnormal{polylog}(n/s)^{-1}s/n$, where $\textnormal{polylog}(n/s)$ is a function that is bounded away from 0 and increases at most polynomially with the log of the inverse sampling ratio $\log(n/s)$.
The proof of Theorem 4.4 is an immediate consequence of Theorem 5 of Athey et al. (2019). Theorems 4.3 and 4.4 demonstrate that the GRF-FPT estimator is able to meet key statistical guarantees.
@@ 5 Applications
where $\nu^{*}(\cdot)$ is a nuisance intercept function, $\theta^{*}(x)=(\theta^{*}_{1}(x),\ldots,\theta^{*}_{K}(x))^{\top}$ are the target effect functions local to $X_{i}=x$, under the assumptions $\mathbb{E}[\epsilon_{i}\mid X_{i}=x]=0$ and $\mathbb{E}[\epsilon_{i}W_{i}\mid X_{i}=x]={\bf 0}$.
Varying coefficient models (VCM). Given regressors $W_{i}\in\mathbb{R}^{K}$, models of the form (28) can be characterized as varying coefficient models (Hastie & Tibshirani, 1993). As discussed in Section 3.2, we must also assume that the regressors $W_{i}$ are conditionally exogenous given $X_{i}=x$.
@@
$\psi_{\theta,\nu}(Y_{i},W_{i})\coloneqq\begin{bmatrix}(Y_{i}-W_{i}^{\top}\theta-\nu)\cdot W_{i}\\ Y_{i}-W_{i}^{\top}\theta-\nu\end{bmatrix}.$
@@
$A_{P}=-\frac{1}{n_{P}}\sum_{\{i:X_{i}\in P\}}(W_{i}-\overline{W}_{P})(W_{i}-\overline{W}_{P})^{\top}.$ (30)
@@
$\hat{\theta}_{P}\coloneqq-A_{P}^{-1}\frac{1}{n_{P}}\sum_{\{i:X_{i}\in P\}}(W_{i}-\overline{W}_{P})(Y_{i}-\overline{Y}_{P}).$ (31)
@@
$\rho^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}_{i}$ $\coloneqq-\xi^{\top}\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(Y_{i},W_{i}),$
@@
The relationship $\rho_{i}^{\texttt{grad}}=A_{P}^{-1}\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ reveals a significant benefit of FPT pseudo-outcomes. The form of $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ eliminates the computational cost associated with the multiplication of $A_{P}^{-1}$, leading to $\mathcal{O}(K^{3})$ computational savings. Furthermore, the computation of $\hat{\theta}_{P}$ in (32) no longer requires solving for $A_{P}^{-1}$. Therefore, we can further enhance computational efficiency by using an accelerated form of pseudo-outcome $\phi_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ instead of $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$:
$\phi_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}\coloneqq-(W_{i}-\overline{W}_{P})\left(Y_{i}-\overline{Y}_{P}-(W_{i}-\overline{W}_{P})^{\top}\tilde{\theta}_{P}\right),$ (33)
where $\hat{\theta}_{P}$ is replaced by $\tilde{\theta}_{P}$ in (32), which is defined as a one-step gradient descent approximation of $\hat{\theta}_{P}$ taken from the origin:
$\tilde{\theta}_{P}\coloneqq\gamma\frac{1}{n_{P}}\sum_{\{i:X_{i}\in P\}}(W_{i}-\overline{W}_{P})(Y_{i}-\overline{Y}_{P}).$ (34)
Here, $\gamma$ denotes the exact line search step size for the regression of $Y_{i}-\overline{Y}_{P}$ on $W_{i}-\overline{W}_{P}$ over $P$:
@@
where $W=\left[\begin{matrix}W_{1}~{}\cdots~{}W_{n_{P}}\end{matrix}\right]^{\top}$ and $Y=\left[\begin{matrix}Y_{1}~{}\cdots~{}Y_{n_{P}}\end{matrix}\right]^{\top}$ with the notation $W-\overline{W}_{P}$ and $Y-\overline{Y}_{P}$ understood as row-wise centering.
The computational cost associated with $\tilde{\theta}_{P}$ is comparatively small because many of the products that appear in (34) and (35) are already computed as part of $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ in (32). Meanwhile, we show in Appendix B.3 that the approximation for the FPT child estimator:
$\bar{\theta}_{C_{j}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}\coloneqq\hat{\theta}_{P}+\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\phi_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}},$
is consistent for the original FPT child estimator $\tilde{\theta}_{C_{j}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ as $\lVert\tilde{\theta}_{C_{j}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}-\bar{\theta}_{C_{j}}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}\rVert=o_{P}(1)$, meaning that this approximation does not alter the asymptotic behavior of our estimator. These accelerations are particularly compelling when the dimension of $\theta^{*}(x)$ is large and computational efficiency is critical, as in large-scale A/B testing with multiple concurrent treatment arms or observational studies with numerous treatment levels (Kohavi et al., 2013; Bakshy et al., 2014).
@@ 6 Simulations
In this section, we perform empirical evaluations of the computational efficiency and estimation accuracy of the GRF-FPT method. We let GRF-FPT1 denote the FPT algorithm using the exact form of the FPT VCM/HTE pseudo-outcomes (32) and we let GRF-FPT2 denote the accelerated FPT algorithm based on the form of the FPT pseudo-outcome approximation (33) in Section 5. We compare both implementations relative to GRF-grad under VCM and HTE designs. Implementation details and links to the reproducible code are found in Appendix C.4.
Settings. We follow the structural model in (28). The auxiliary variables $X_{i}$ are drawn from the Gaussian copula with latent covariance matrix $\Sigma$, where $[\Sigma]_{j,k}=(0.3)^{|j-k|}$. Supporting experiments for multicollinearity in $X_{i}$ can be found in Appendix D.2. The outcomes $Y_{i}$ follow (28) with Gaussian noise $\epsilon_{i}\sim\mathcal{N}(0,1)$. For VCM experiments, regressors $W_{i}\in\mathbb{R}^{K}$ are sampled from $\mathcal{N}_{K}({\bm{0}},\mathbb{I})$. For HTE experiments, $W_{i}\in\{0,1\}^{K}$ follows a multinomial distribution, $W_{i}\mid X_{i}=x\sim\text{Multinomial}(1,(\pi_{1}(x),\ldots,\pi_{K}(x)))$, where $\pi_{k}(x)$ is the probability of treatment level $k\in\{1,\ldots,K\}$, characterizing a variety of different location-specific dependence structures through the setting of $\pi_{k}(\cdot)$. We set $\nu^{*}(x)\coloneqq 0$ and vary the target effect functions $\theta^{*}_{k}(x)$ and treatment probabilities $\pi_{k}(x)$ across different settings, fully detailed in Appendix C.4. Throughout our experiments we use subsampling ratio $s/n=0.5$. Supporting experiments under different subsample ratios are found in Appendix D.2.
Results. The relative computational advantage of forests trained under GRF-FPT is displayed in Figure 3, while Figure 5 (in Appendix D.3) summarizes the absolute fit times across the three methods. These data show that the FPT mechanism is able to consistently offer a relative advantage, observing speedups of up to 3.5$\times$ faster than the gradient-based approach at the largest dimension $K=256$. Figure 3 also shows increasing gains with increasing $K$ and provides an empirical measurement of the theoretical scaling benefits discussed in Section 5. Moreover, the absolute fit times in Figure 5 (in Appendix D.3) illustrate that our method consistently remains faster than GRF-grad, with no clear computational or algorithmic bottleneck as a function of either $n$ or $K$. Supporting experiments exploring the effects of sample sizes up to $n=500,000$ are presented in Appendix D.2, while Figures 7 and 8 (in Appendix D.3) show that even when $n$ is small, GRF-FPT still observes a noticeable gain relative to GRF-grad. Additional timing benchmarks for VCM experiments and all HTE experiments are discussed in Appendix D.3.
To assess estimation accuracy, we evaluate the mean squared error (MSE) of $\hat{\theta}(x)$ across 50 replications of the model and testing on a separate set of $5,000$ observations. Figure 6 in Appendix D.3 confirms that GRF-FPT matches the accuracy of GRF-grad, while significantly reducing computation time. Further comparisons for both VCM and HTE settings are provided in Appendix D.3.
@@ 7 Real Data Application
Data. In this section we apply GRF-FPT to the analysis of geographically-varying effects $\theta^{*}(x)$ on housing prices. The data, first appearing in Kelley Pace & Barry (1997), contains 20,640 observations of housing prices taken from the 1990 California census. Each observation corresponds to measurements aggregated over a small geographical census block, and contains measurements of 9 variables: median housing value, longitude, latitude, median housing age, total rooms, total bedrooms, population, households, and median income. We employ a VCM design of the form (28) where $Y_{i}$ denotes the housing value, $X_{i}$ denote the spatial coordinates, and $W_{i}=(W_{i,1},\ldots,W_{i,6})^{\top}$ are the remaining six regressors. Details of the model and data transformations used for the California housing analysis is found in Appendix F.
Results. Table 7 summarizes the computational benefit of GRF-FPT applied to the California housing data. Figure 4 illustrates the six geographically-varying effect estimates under GRF-FPT2, with qualitatively similar results shown in Figure 16 for GRF-FPT1 and GRF-grad in Appendix F. Figure 4 shows clearly the geographically-dependent relationship between different housing features and housing prices. In major urban centers such as LA, San Francisco, and Sacramento, housing prices tend to decrease with an increasing number of households, and may reflect overcrowding in densely populated areas. In contrast, rural regions show the opposite trend: prices rise slightly when rural areas have a larger number of housing units. This suggests that, in sparsely populated rural areas, a modest increase in households makes these places more attractive and livable. Median income, however, consistently shows a positive effect on prices across nearly all of California, while population size tends to show a negative effect, highlighting broader state-wide pressures on housing affordability.
Figure 4: Geographically-varying GRF-FPT2 estimates $\hat{\theta}(x)$.
@@ 8 Conclusion
Our results demonstrate that the FPT algorithm offers a substantial computational advantage over GRF-grad with comparable statistical accuracy, and highlights GRF-FPT as a powerful method for multi-dimensional estimation, particularly when estimates of the target function must be learned from the data rather than observed directly. Future work may explore extensions to larger-scale problems and alternative estimation tasks, as in unsupervised learning and structured prediction. Our findings position GRF-FPT as a scalable and robust alternative for practitioners seeking efficient localized estimation.
Impact Statement
@@
Acknowledgments
This work was supported by Natural Sciences and Engineering Research Council (NSERC) Discovery Grant (RGPIN-2024-06780) and FRQNT Team Research Project Grant (FRQ-NT 327788).
@@ A.1 Assumptions
We follow the key assumptions of Athey et al. (2019) made for the theoretical analyses of GRF. The predictor and parameter spaces are both subsets of Euclidean space such that $x\in\mathcal{X}=[0,1]^{p}$ and $(\theta,\nu)\in\mathcal{B}\subset\mathbb{R}^{K}$, where $\mathcal{B}$ is a compact subset of $\mathbb{R}^{K}$. Under the analyses of Wager & Walther (2015), we suppose that the features of the auxiliary covariates $X_{i}=(X_{i,1},\ldots,X_{i,p})^{\top}$ have density $f_{X}$ that is bounded away from 0 and $\infty$, i.e. $c\leq f_{X}(x)\leq C<\infty$, for some constants $c>0$ and $C<\infty$. GRF does not require that the score function $\psi$ is continuous in $(\theta,\nu)$, as is the case for quantile estimation, one does require that the expected score/moment function
$M_{\theta,\nu}(x)\coloneqq\mathbb{E}_{O|X}\left[\psi_{\theta,\nu}(O)\mid X=x\right],$ (36)
is smoothly varying in its parameters $(\theta,\nu)$.
@@
    $V(x)=\begin{bmatrix}V_{\theta\theta}(x)&V_{\theta\nu}(x)\\ V_{\nu\theta}(x)&V_{\nu\nu}(x)\end{bmatrix}.$ (37)
@@
    The score functions $\psi_{\theta,\nu}(O_{i})$ have a continuous covariance structure in the following sense: Let $\gamma(\cdot,\cdot)$ denote the worst-case variogram:
@@
    The score function $\psi_{\theta,\nu}(O_{i})$ can be written as
@@
    where $\lambda$ is Lipschitz-continuous in $(\theta,\nu)$, $g:\{O_{i}\}\to\mathbb{R}$ a univariate summary of the observables $O_{i}$, and $\zeta_{\theta}:\mathbb{R}\to\mathbb{R}$ any family of monotone and bounded functions.
@@
    For any weights $\alpha_{i}$ with $\sum\alpha_{i}=1$, the minimizer $(\hat{\theta},\hat{\nu})$ of the weighted empirical estimation problem (5) satisfies:
@@
    The score function $\psi_{\theta,\nu}(O_{i})$ is a negative subgradient of a convex function, and the moment function $M_{\theta,\nu}(X_{i})$ is the negative gradient of a strongly convex function.